# Data-Driven Distributed Output Synchronization of Heterogeneous Discrete-Time Multi-Agent Systems

### v1

@@ V Conclusions
Since $\Psi$ is of full row rank, its right inverses are those and those only that can be expressed as
$\Psi^{\#}=\Psi^{\dagger}+\left(I_{r}-\Psi^{\dagger}\Psi\right)Q,$ (37)
as $Q$ varies in ${\mathbb{R}}^{r\times(n+k)}$. Therefore if $C$ is expressed as in (36), for some $Q$, then obviously (35) holds. Conversely, if $C$ satisfies (35), the fact that $B$ is of full row rank ensures that
$\begin{bmatrix}A\cr B\end{bmatrix}\begin{bmatrix}C&B^{\dagger}\end{bmatrix}=\begin{bmatrix}I_{n}&*\cr{\mathbb{0}}&I_{k}\end{bmatrix},$
for some matrix * and hence there exists $B^{\#}$ such that
$\Psi\begin{bmatrix}C&B^{\#}\end{bmatrix}=\begin{bmatrix}A\cr B\end{bmatrix}\begin{bmatrix}C&B^{\#}\end{bmatrix}=\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&I_{k}\end{bmatrix}.$
This immediately implies that $\begin{bmatrix}C&B^{\#}\end{bmatrix}$ is a right inverse of $\Psi$ and hence can be expressed as in (37) for some $Q$. Therefore (36) holds.

### v2

@@ Appendix A Appendix
Let $L$ and $R$ be a full column rank matrix and a full row rank matrix, respectively, so that $B=LR$. Then
$\bar{\Psi}=\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&L\end{bmatrix}\begin{bmatrix}A\cr R\end{bmatrix}$.
Note that in the previous factorization the matrix on the left is of full column rank, while the matrix on the right is of full row rank and such common rank coincides with ${\rm rank}(\bar{\Psi})$. Consequently, by the properties of the Moore-Penrose inverse [10], we can claim that
$\bar{\Psi}^{\dagger}=\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&L\end{bmatrix}^{\dagger}$.
Clearly, if (36) holds for some $Q$, then
$\bar{\Psi}C$ $=\bar{\Psi}\left[\bar{\Psi}^{\dagger}+\left(I_{r}-\bar{\Psi}^{\dagger}\bar{\Psi}\right)Q\right]\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$
$=\bar{\Psi}\bar{\Psi}^{\dagger}\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}=\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&LL^{\dagger}\end{bmatrix}\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}=\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$,
which means that (35) holds.
Conversely, suppose that (35) holds. Since $L$ is of full column rank, this implies that
$\begin{bmatrix}A\cr R\end{bmatrix}C=\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$.
Since $R$ is of full row rank, its Moore-Penrose inverse $R^{\dagger}$ is also a right inverse, i.e., $RR^{\dagger}=I_{\bar{k}}$, and hence
$\begin{bmatrix}A\cr R\end{bmatrix}\begin{bmatrix}C&R^{\dagger}\end{bmatrix}=\begin{bmatrix}I_{n}&AR^{\dagger}\cr{\mathbb{0}}&I_{\bar{k}}\end{bmatrix}$.
Consequently, upon setting $R^{\#}\doteq R^{\dagger}-CAR^{\dagger}$, we get
$\begin{bmatrix}A\cr R\end{bmatrix}\begin{bmatrix}C&R^{\#}\end{bmatrix}=\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&I_{\bar{k}}\end{bmatrix}$.
Since $\begin{bmatrix}C&R^{\#}\end{bmatrix}$ is a right inverse of $\begin{bmatrix}A\cr R\end{bmatrix}$, there exists some matrix $Q$ such that
$\begin{bmatrix}C&R^{\#}\end{bmatrix}=\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}+\left(I_{r}-\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}A\cr R\end{bmatrix}\right)Q.$
This implies that
$C=\left(\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}+\left(I_{r}-\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}A\cr R\end{bmatrix}\right)Q\right)\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$, (37)
for some matrix $Q$. It is easy to see that
$\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}=\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&L\end{bmatrix}^{\dagger}\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}=\bar{\Psi}^{\dagger}\begin{bmatrix}I_{n}\cr{\mathbb{0}}\end{bmatrix}$
as well as
$\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}A\cr R\end{bmatrix}=\begin{bmatrix}A\cr R\end{bmatrix}^{\dagger}\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&L\end{bmatrix}^{\dagger}\begin{bmatrix}I_{n}&{\mathbb{0}}\cr{\mathbb{0}}&L\end{bmatrix}\begin{bmatrix}A\cr R\end{bmatrix}=\bar{\Psi}^{\dagger}\bar{\Psi}.$
This shows that (37) coincides with (36), and hence $C$ can be written as in (36) for some $Q$. ∎